# Human-Activity AGV Quality Assessment: A Benchmark Dataset and an Objective Evaluation Metric

### v1

@@ Abstract
AI-driven video generation techniques have made significant progress in recent years. However, AI-generated videos (AGVs) involving human activities often exhibit substantial visual and semantic distortions, hindering the practical application of video generation technologies in real-world scenarios. To address this challenge, we conduct a pioneering study on human activity AGV quality assessment, focusing on visual quality evaluation and the identification of semantic distortions. First, we construct the AI-Generated Human activity Video Quality Assessment (Human-AGVQA) dataset, consisting of $3,200$ AGVs derived from $8$ popular text-to-video (T2V) models using $400$ text prompts that describe diverse human activities. We conduct a subjective study to evaluate the human appearance quality, action continuity quality, and overall video quality of AGVs, and identify semantic issues of human body parts. Based on Human-AGVQA, we benchmark the performance of T2V models and analyze their strengths and weaknesses in generating different categories of human activities. Second, we develop an objective evaluation metric, named AI-Generated Human activity Video Quality metric (GHVQ), to automatically analyze the quality of human activity AGVs. GHVQ systematically extracts human-focused quality features, AI-generated content-aware quality features, and temporal continuity features, making it a comprehensive and explainable quality metric for human activity AGVs. The extensive experimental results show that GHVQ outperforms existing quality metrics on the Human-AGVQA dataset by a large margin, demonstrating its efficacy in assessing the quality of human activity AGVs. The Human-AGVQA dataset and GHVQ metric will be released in public at GitHub.
1 Introduction
@@
Figure 1: Overview of this work. This study aims to address the distortions in human activity AGVs, with some typical distortions shown in the first row. To achieve this goal, we introduce the Human-AGVQA dataset, consisting of $3,200$ human activity AGVs with the quality labels, and the GHVQ metric to objectively compute the quality of human activity AGVs.
@@
AI-driven video generation, particularly in the form of text-to-video (T2V) generation [9, 1], can automatically create visual realism videos according to text descriptions. This convenient and cost-effective method of video creation has significant potential for applications in entertainment, art, advertising, education, and various other fields. However, many studies [33, 53] indicate that current T2V models still struggle to generate realistic human figures and actions. For instance, AI-generated videos (AGVs) may exhibit distorted, incomplete, or abnormal body parts and actions, as illustrated in Figure 1. These artifacts can significantly affect the user’s perceptual quality, as the human body is highly structured, and the human visual system is susceptible to irregular body form and movement irregularities.
@@
Therefore, accurately assessing the quality of human-activity AGVs is critical to automatically monitor the visual quality of large-scale AGVs in video generation applications, measure the progress of T2V models, and serve as an optimization or reward function to enhance the capability of T2V models. Unfortunately, the quality assessment for AI-generated content (AIGC) [54] is still in its infancy. The general image/video quality assessment (I/VQA) metrics have been shown to perform poorly in evaluating AGVs [84, 39], while common used metrics in T2V studies, such as Inception Score (IS) [65], Fréchet Inception Distance (FID) [28], and Fréchet Video Distance (FVD) [79], evaluate the generation performance and diversity of T2V models by comparing a set of AGVs to real data distributions in feature Inception feature spaces, thus failing to reflect the quality at the individual video level.
@@
To bridge this gap, we establish the first quality assessment benchmark for human activity AGVs, providing a comprehensive analysis of quality issues in human activity videos generated by mainstream T2V models and serving as a testbed for validating the effectiveness of quality assessment metrics in evaluating human activity AGVs. Specifically, we first construct the AI-Generated Human activity Video Quality Assessment (Human-AGVQA) dataset, which contains $3,200$ AGVs derived by $8$ state-of-the-art T2V models using $400$ text prompts. We define two practical quality assessment problems for human activity AGVs. The first is quality scoring, which quantifies the specific quality score of AGVs, allowing us to compare the quality of one AGV or T2V model against others. The second one is semantic issue identification, which pinpoints problematic body parts that affect quality scores, providing insights for further optimization of T2V models. Therefore, for the quality scoring task, we invite $80$ subjects to rate the quality scores of AGVs based on three critical dimensions: human appearance quality, action continuity quality, and overall visual quality. For the semantic issue identification task, we invite $5$ experts to label whether the human bodies (face, arms, torso, legs, and feet) exhibit semantic issues. Based on these subjectively rated quality labels, we can benchmark the performance of T2V models and analyze their strengths and weaknesses in generating different categories of human activities.
@@
  • We establish the Human-AGVQA dataset for assessing the quality of human activity AGVs. Human-AGVQA comprises $3,200$ videos generated by $8$ T2V models using $400$ text prompts, along with subjectively rated quality labels for $2$ types of assessment tasks. The diversity of text prompts and the richness of quality labels make our benchmark well-suited for investigating the quality assessment of human activity AGVs.
@@
  • We propose the GHVQ metric to automatically evaluate the human appearance quality, action continuity quality, and overall video quality of AGVs, and identify whether five human body parts exhibit semantic issues. Experimental results indicate that GHVQ significantly outperforms existing related quality metrics, demonstrating its effectiveness.
2 Related Works
@@
Video Generation Techniques can be broadly categorized into three types: GAN/VAE-based methods [59, 62, 46, 14], autoregressive-based methods [86, 49, 31, 80, 3, 20], and diffusion-based methods [30, 25, 90, 104, 100, 8, 7, 67, 29, 35, 94]. Among these, diffusion-based methods have seen significant progress, with some commercial models [16, 1] already being developed or launched for real-world applications. However, several studies [70, 13, 53, 33] point out that current video generation models still struggle with generating realistic human characteristics and actions, which emphasizes the need for quality assessment study in human activity AGVs.
Video Quality Assessment studies can be divided into two categories: knowledge-driven methods and data-driven methods. Knowledge-driven BVQA methods leverage prior knowledge to extract quality features (e.g., natural scenes statistics features) for quality regression, such as TLVQM [38], VIDEVAL [77], RAPIQUE [78], etc. In contrast, data-driven BVQA methods automatically learn the quality-aware features by training a carefully designed deep neural network (DNN) in a learning-based manner. Popular data-driven methods generally 1) use a 2D network to extract key frame features and a 3D network to capture motion features from video chunks, as seen in models like BVQA [41], SimpleVQA [71], PatchVQA [96], MinimalisticVQA [73], etc., or 2) employ a 3D network directly to extract video-level features from video chunks, such as FastVQA [87]. The 2D and 3D backbones can be pre-trained on other computer vision tasks [95, 24] or fine-tuned on VQA datasets in an end-to-end manner. Since current VQA research primarily focuses on natural videos, these methods often show limited performance when assessing the quality of AGVs [52].
@@
Figure 2: Illustration of the MOSs distribution. HA, AC, and OV quality represent human appearance, action continuity, and overall video, respectively. These abbreviations are used consistently throughout the paper.
@@
AGV Quality Assessment has been explored in several studies [54, 33, 39]. For instance, FETV [53] benchmarks five T2V models based on static quality, temporal quality, overall alignment, and fine-grained alignment. EvalCrafter [54] evaluates 17 objective metrics, including visual quality, text-video alignment, motion quality, and temporal consistency, on a subjectively-rated AGV dataset. It also proposes a combined metric that leads to improved performance. VBench [33] introduces 16 additional dimensional metrics to evaluate the performance of T2V models. Moreover, [39] constructs the T2VQA-DB, a large-scale text-to-video quality assessment dataset containing 10,000 videos, and develops the T2VQA metric, which focuses on the overall quality of AGVs. Despite these advancements, no specific metric has been developed to evaluate the quality of human activity AGVs.
3 Human-AGVQA Dataset
3.1 Human activity AGVs Collection
@@
Text Prompts Selection. Text prompts describe the video content generated by T2V models. To comprehensively analyze the quality assessment problem of human activity AGVs, the selected text prompts in our dataset should cover a wide range of real-world human activities. Therefore, we classify the properties of these words into ages, genders, ethnicity, occupations, scenes, emotion, appearance, and activities categories and further divide each category into $44$ subcategories that frequently occur in daily human life. We then randomly select one subcategory (e.g., sports) from the activity category, one subcategory (e.g., outdoor) from the scene category, and two or three subcategories from human appearance-based categories (e.g., athlete from the occupation category and male from the gender category). These keywords are combined into a text prompt using GPT-4 [61] with a prompt:
@@
# Please construct a complete sentence based on the following four concepts: “sports”, “outdoor”, “athlete”, and “male”. Note that the specific words “sports”, “outdoor”, “athlete”, and “male” are not required to appear in the sentence, but the sentence should clearly convey the ideas related to these four categories. #.
@@
Figure 3: The MOS comparison for $8$ T2V models.
T2V Models Selection. We select eight state-of-the-art (SOTA) T2V models, including Gen-2 [16], StableVideo [1], StreamingT2V [26], Video Crafter2 [9], MagicTime [98], Open Sora [40], Latte [55], and AnimateDiff [22], to generate the video for each prompt. We summary the detailed information about these T2V models in the supplementary material.
Figure 4: The MOS comparison for $8$ text categories and their $44$ subcategories.
In summary, a total of $3,200$ videos were generated by $8$ SOTA T2V models using $400$ diverse text prompts in the Human-AGVQA dataset.
3.2 Quality Assessment Dimensions
@@
Semantic Distortion Identification. Most quality metrics only provide a specific numerical value to quantify video quality, and cannot indicate which parts of the video are distorted or need further enhancement or optimization. This limitation restricts their applicability in quality artifact analysis and fine-grained video quality optimization. To address this, we label the semantic distorted human body parts in AGVs, providing a detailed explanation of quality scores. Specifically, we first annotate whether five body parts (face, arms  [^1] Here, the arms include the hands., torso, legs, feet) are present in AGVs. Then, for each detected body part, we label whether it contains semantic artifacts. Finally, we provide $10$ binary labels per AGV for semantic distortion identification.
3.3 Subjective Quality Experiment
A total of $80$ subjects participated in the visual quality scoring experiment, with ages ranging from $20$ to $30$ years. The group included $46$ males and $34$ females. Given that the semantic artifact identification task is less complex than the visual quality scoring task, we invited $5$ experts in the field of AIGC quality assessment to perform the semantic artifact identification. $3,200$ videos were divided into eight groups, each containing $400$ videos that covered all 400 prompts, and each video was rated by $10$ subjects, and labeled by $5$ experts. In total, there are $96,000$ opinion scores and $160,000$ binary labels in the Human-AGVQA dataset. Additional details on the subjective experimental settings can be found in the supplemental material.
3.4 Data Processing and Analysis
For the quality scoring task, we follow the recommended method in  [2] to process the raw subjective ratings into the mean opinion scores (MOSs). For the distortion identification task, we use a voting method to determine body presence and body distortion. The label with the most votes is selected as the final result. The details about data processing and inter-subject consistency are shown in supplementary material.
3.4.1 MOS Distribution Analysis
@@
Figure 5: The percentage of AGVs that contain human body parts (a) and the percentage of these body parts exhibit semantic distortions (b).
Figure 6: The GHVQ framework comprises four essential modules: a spatial quality analyzer for extracting human-focused and holistic quality features, an action quality analyzer for capturing temporal continuity of the body movement, a text feature extractor for identifying video content that needs to be generated by text prompts, and a quality regressor for mapping these quality-aware features to multi-dimensional quality scores and semantic distortion labels.
3.4.2 MOS Analysis for T2V Models
We calculate the average MOSs of eight T2V models across three quality dimensions to assess their strengths and weaknesses in generating human activity videos. The results are shown in Figure 3. Gen-2 and StableVideo demonstrate superior performance compared to the other T2V models, especially in human appearance quality and overall video quality. Notably, both Gen-2 and StableVideo are commercial T2V tools, indicating that the generation quality of commercial algorithms continues to surpass that of academic open-source models. AnimateDiff and VideoCrafter2 show similar performance to Gen-2 and StableVideo in terms of action continuity quality, highlighting their temporal consistency modeling capabilities.
3.4.3 MOS Analysis for Text Prompts

### v2

@@ Abstract.
AI-driven video generation techniques have made significant progress in recent years. However, AI-generated videos (AGVs) involving human activities often exhibit substantial visual and semantic distortions, hindering the practical application of video generation technologies in real-world scenarios. To address this challenge, we conduct a pioneering study on human activity AGV quality assessment, focusing on visual quality evaluation and the identification of semantic distortions. First, we construct the AI-Generated Human activity Video Quality Assessment (Human-AGVQA) dataset, consisting of $6,000$ AGVs derived from $15$ popular text-to-video (T2V) models using $400$ text prompts that describe diverse human activities. We conduct a subjective study to evaluate the human appearance quality, action continuity quality, and overall video quality of AGVs, and identify semantic issues of human body parts. Based on Human-AGVQA, we benchmark the performance of T2V models and analyze their strengths and weaknesses in generating different categories of human activities. Second, we develop an objective evaluation metric, named AI-Generated Human activity Video Quality metric (GHVQ), to automatically analyze the quality of human activity AGVs. GHVQ systematically extracts human-focused quality features, AI-generated content-aware quality features, and temporal continuity features, making it a comprehensive and explainable quality metric for human activity AGVs. The extensive experimental results show that GHVQ outperforms existing quality metrics on the Human-AGVQA dataset by a large margin, demonstrating its efficacy in assessing the quality of human activity AGVs. The Human-AGVQA dataset and GHVQ metric will be released publicly.
video quality assessment, AI-generated video, human activity video
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing HCI design and evaluation methodsccs: Computing methodologies Computer vision problems
@@
Figure 1. Overview of this work. This study aims to address the distortions in human activity AGVs, with some typical distortions shown in the first column. To achieve this goal, we introduce the Human-AGVQA dataset, consisting of $6,000$ human activity AGVs with quality and distortion identification labels, and we benchmark over $40$ quality assessment and MLLM methods on it. Finally, the proposed GHVQ metric objectively computes the quality of human activity AGVs.
1. Introduction
@@
AI-driven video generation, particularly in the form of text-to-video (T2V) generation (Chen et al., 2024a; AI, 2024a), can automatically create visual realism videos according to text descriptions. This convenient and cost-effective method of video creation has significant potential for applications in entertainment, art, advertising, education, and various other fields. However, many studies (Huang et al., 2024; Liu et al., 2023) indicate that current T2V models still struggle to generate realistic human figures and actions. For instance, AI-generated videos (AGVs) may exhibit distorted, incomplete, or abnormal body parts and actions, as illustrated in Figure 1. These artifacts can significantly affect the user’s perceptual quality, as the human body is highly structured, and the human visual system is susceptible to abnormal body parts and irregular movement.
@@
Therefore, accurately assessing the quality of human-activity AGVs is critical to automatically monitor the visual quality of large-scale AGVs in video generation applications, measure the progress of T2V models, and serve as an optimization or reward function to enhance the capability of T2V models. Unfortunately, the quality assessment for AI-generated content (AIGC) (Liu et al., 2024a) is still in its infancy. The general image/video quality assessment (I/VQA) metrics have been shown to perform poorly in evaluating AGVs (Wang et al., 2024b; Kou et al., 2024), while common used metrics in T2V studies, such as Inception Score (IS) (Salimans et al., 2016), Fréchet Inception Distance (FID) (Heusel et al., 2017), and Fréchet Video Distance (FVD) (Unterthiner et al., 2018), evaluate the generation performance and diversity of T2V models by comparing a set of AGVs to real data distributions in feature Inception feature spaces, thus failing to reflect the quality at the individual video level.
@@
To bridge this gap, we establish the first quality assessment benchmark for human activity AGVs, providing a comprehensive analysis of quality issues in human activity videos generated by mainstream T2V models and serving as a testbed for validating the effectiveness of quality assessment metrics in evaluating human activity AGVs. Specifically, we first construct the AI-Generated Human activity Video Quality Assessment (Human-AGVQA) dataset, which contains $6,000$ AGVs derived by $15$ state-of-the-art T2V models using $400$ text prompts. We define two practical quality assessment problems for human activity AGVs. The first is quality scoring, which quantifies the specific quality score of AGVs, allowing us to compare the quality of one AGV or T2V model against others. The second one is semantic issue identification, which pinpoints problematic body parts that affect quality scores, providing insights for further optimization of T2V models. Therefore, for the quality scoring task, we invite $80$ subjects to rate the quality scores of AGVs based on three critical dimensions: human appearance quality, action continuity quality, and overall visual quality. For the semantic issue identification task, we invite $5$ experts to label whether the human bodies (face, hands, arms, torso, legs, and feet) exhibit semantic issues. Based on these subjectively rated quality labels, we can benchmark the performance of T2V models and analyze their strengths and weaknesses in generating different categories of human activities.
@@
  • We establish the Human-AGVQA dataset for assessing the quality of human activity AGVs. Human-AGVQA comprises $6,000$ videos generated by $15$ T2V models using $400$ text prompts, along with subjectively rated quality labels for $2$ types of assessment tasks. The diversity of text prompts and the richness of quality labels make our benchmark well-suited for investigating the quality assessment of human activity AGVs.
@@
  • We propose the GHVQ metric to automatically evaluate the human appearance quality, action continuity quality, and overall video quality of AGVs, and identify whether six human body parts exhibit semantic issues. Experimental results indicate that GHVQ significantly outperforms existing related quality metrics, demonstrating its effectiveness.
2. Related Works
@@
Video Generation Techniques can be broadly categorized into three types: GAN/VAE-based methods (Pan et al., 2017; Li et al., 2018), autoregressive-based methods (Liang et al., 2022; Hong et al., 2022; Villegas et al., 2022), and diffusion-based methods (Ho et al., 2022; Zeng et al., 2023; Chen et al., 2024b). Among these, diffusion-based methods have seen significant progress, with some commercial models (Esser et al., 2023; AI, 2024a) already being developed or launched for real-world applications. However, several studies (Sun et al., [n. d.]; Cho et al., 2024; Liu et al., 2023; Huang et al., 2024) point out that current video generation models still struggle with generating realistic human characteristics and actions, which emphasizes the need for quality assessment study in human activity AGVs.
@@
Figure 2. Illustration of the MOSs distribution. HA, AC, and OV quality represent human appearance, action continuity, and overall video, respectively. These abbreviations are used consistently throughout the paper.
Video Quality Assessment studies can be divided into two categories: knowledge-driven methods and data-driven methods. Knowledge-driven VQA methods leverage prior knowledge to extract quality features (e.g., natural scenes statistics features) for quality regression, such as TLVQM (Korhonen, 2019), VIDEVAL (Tu et al., 2021a), RAPIQUE (Tu et al., 2021b), etc. In contrast, data-driven BVQA methods automatically learn the quality-aware features by training a carefully designed deep neural network (DNN). Popular data-driven VQA methods generally 1) use a 2D network to extract key frame features and a 3D network to capture motion features from video chunks, as seen in models like Li22 (Li et al., 2022b), SimpleVQA (Sun et al., 2022), PatchVQA (Ying et al., 2021), MinimalisticVQA (Sun et al., 2024), etc., or 2) employ a 3D network directly to extract video-level features from video chunks, such as FastVQA (Wu et al., 2022). The 2D and 3D backbones can be pre-trained on other computer vision tasks (Ying et al., 2020; He et al., 2016) or fine-tuned on VQA datasets in an end-to-end manner. Since current VQA research primarily focuses on natural videos, these methods often show limited performance when assessing the quality of AGVs (Liu et al., 2024b).
@@
AGV Quality Assessment has been explored in several studies (Liu et al., 2024a; Huang et al., 2024; Kou et al., 2024). For instance, FETV (Liu et al., 2023) benchmarks five T2V models based on static quality, temporal quality, overall alignment, and fine-grained alignment. EvalCrafter (Liu et al., 2024a) evaluates 17 objective metrics, including visual quality, text-video alignment, motion quality, and temporal consistency, on a subjectively-rated AGV dataset. It also proposes a combined metric that leads to improved performance. VBench (Huang et al., 2024) introduces 16 additional dimensional metrics to evaluate the performance of T2V models. Moreover, (Kou et al., 2024) constructs the T2VQA-DB, a large-scale text-to-video quality assessment dataset containing 10,000 videos, and develops the T2VQA metric, which focuses on the overall quality of AGVs. Despite these advancements, no specific metric has been developed to evaluate the quality of human activity AGVs.
Figure 3. The MOS comparison for $15$ T2V models.
3. Human-AGVQA Dataset
3.1. Human activity AGVs Collection
Figure 4. The MOS comparison for $15$ text categories and their $44$ subcategories.
@@ 3.1. Human activity AGVs Collection
Text Prompts Selection. Text prompts describe the video content generated by T2V models. To comprehensively analyze the quality assessment problem of human activity AGVs, the selected text prompts in our dataset should cover a wide range of real-world human activities. Therefore, we classify the properties of these words into ages, genders, ethnicity, occupations, scenes, emotion, appearance, and activities categories and further divide each category into $44$ subcategories that frequently occur in daily human life. We then randomly select one subcategory (e.g., sports) from the activity category, one subcategory (e.g., outdoor) from the scene category, and two or three subcategories from human appearance-based categories (e.g., athlete from the occupation category and male from the gender category). These keywords are combined into a text prompt using GPT-4 (OpenAI, 2024a) with a prompt:
@@
# Please construct a complete sentence based on the following four concepts: ”sports”, ”outdoor”, ”athlete”, and ”male”. Note that the specific words ”sports”, ”outdoor”, ”athlete”, and ”male” are not required to appear in the sentence, but the sentence should clearly convey the ideas related to these four categories. #.
@@
T2V Models Selection. We select $15$ state-of-the-art (SOTA) T2V models, including Gen-2 (Esser et al., 2023), StableVideo (AI, 2024a), StreamingT2V (Henschel et al., 2024), Video Crafter2 (Chen et al., 2024a), MagicTime (Yuan et al., 2024), Open Sora (Lab and etc., 2024), Latte (Ma et al., 2024), and AnimateDiff (Guo et al., 2024), Sora(OpenAI, 2024b) , Dreamina(ByteDance, 2024), Kling(Technology, 2024), Ying(AI, 2024b), Wan(Cloud, 2025), LTX-Video(HaCohen et al., 2024), Show-1(Zhang et al., 2023a)to generate the video for each prompt. We summary the detailed information about these T2V models in the supplementary material.
In summary, a total of $6,000$ videos were generated by $15$ SOTA T2V models using $400$ diverse text prompts in the Human-AGVQA dataset.
3.2. Quality Assessment Dimensions
@@
Figure 5. The percentage of AGVs that contain human body parts (a) and the percentage of these body parts exhibit semantic distortions (b).
Semantic Distortion Identification. Most quality metrics only provide a specific numerical value to quantify video quality, and cannot indicate which parts of the video are distorted or need further enhancement or optimization. This limitation restricts their applicability in quality artifact analysis and fine-grained video quality optimization. To address this, we label the semantic distorted human body parts in AGVs, providing a detailed explanation of quality scores. Specifically, we first annotate whether $6$ body parts (face, hands, arms, torso, legs, feet) are present in AGVs. Then, for each detected body part, we label whether it contains semantic artifacts. Finally, we provide $12$ binary labels per AGV for semantic distortion identification.
3.3. Subjective Quality Experiment
A total of $80$ subjects participated in the visual quality scoring experiment, with ages ranging from $20$ to $30$ years. The group included $46$ males and $34$ females. Given that the semantic artifact identification task is less complex than the visual quality scoring task, we invited $5$ experts in the field of AIGC quality assessment to perform the semantic artifact identification. $6,000$ videos were divided into $15$ groups, each containing $400$ videos that covered all 400 prompts, and each video was rated by $15$ subjects, and labeled by $5$ experts. In total, there are $270,000$ opinion scores and $360,000$ binary labels in the Human-AGVQA dataset. Additional details on the subjective experimental settings can be found in the supplemental material.
3.4. Data Processing and Analysis
For the quality scoring task, we follow the recommended method in  (Met, 2002) to process the raw subjective ratings into the mean opinion scores (MOSs). For the distortion identification task, we use a voting method to determine body presence and body distortion. The label with the most votes is selected as the final result. The details about data processing and inter-subject consistency are shown in supplementary material.
3.4.1. MOS Distribution Analysis
@@
3.4.2. MOS Analysis for T2V Models
We calculate the average MOSs of $15$ T2V models across three quality dimensions to assess their strengths and weaknesses in generating human activity videos. The results are shown in Figure 3. Kling, Sora, and Wan are capable of producing highly realistic human appearances and action continuity with excellent overall quality, demonstrating strong generation quality. Dreamina, Ying, StableVideo, and Gen-2 also show strong performance. Notably, StableVideo and Gen2 exhibit significantly higher OV quality compared to the other two dimensions. These above models are well-known commercial T2V tools, reinforcing the trend that commercial algorithms often surpass open-source models in terms of output quality. LTX-Video, Show-1, AnimateDiff, and VideoCrafter2 show moderate performance across the dimensions, indicating that they have room for improvement. Latte, MagicTime, OpenSora, and StreamingT2V have lower scores across all three quality dimensions, which suggests that these models may face challenges in generating high-quality videos.
3.4.3. MOS Analysis for Text Prompts